# Mykola Khrypchenko

### v3

@@ 3.2 Algebras with enough idempotents
$am=\prod_{i,j\in I}\left(\sum_{k\in I}a_{ik}m_{kj}\right)\text{ and }ma=\prod_{i,j\in I}\left(\sum_{k\in I}m_{ik}a_{kj}\right)$ (13)
@@
The operations 13 are clearly well-defined and bilinear. For all $a,b\in A$ and $m\in M$ we have
@@
Since $I_{1}:=\{i\in I\mid\exists k\in I:\ a_{ik}\neq 0\}$ and $I_{2}:=\{j\in I\mid\exists l\in I:\ b_{lj}\neq 0\}$ are finite, then $\operatorname{supp}(amb)$ is finite as a subset of $I_{1}\times I_{2}$ by 14. Thus, $AMA\subseteq A$. The converse inclusion is obvious because $a_{i}=e_{i}\cdot e_{i}\cdot a_{i}$ for all $a_{i}\in e_{i}A$, so that $A=A\cdot A\cdot A\subseteq AMA$. ∎
@@
The fact that any $\mathrm{id}$-matching with $\cdot$ product * on $A$ has form 16 is a consequence of Lemma 3.17. Conversely, let * be given by 16. Then
@@
Thus, 17 holds for all $a\in A$.
Conversely, assume 17. Then
@@
We have $e_{k}\cdot(e_{i}*a_{ij})=(e_{k}*e_{i})\cdot a_{ij}=0$ for all $a_{ij}\in e_{i}Ae_{j}$ and $k\neq i$ by Lemma 3.20, so that
@@
But $e_{i}*a_{ij}=e_{i}*(a_{ij}\cdot e_{j})=(e_{i}\cdot a_{ij})*e_{j}=a_{ij}*e_{j}$, whence 19 by 20 and 21. ∎
@@
Let $\varphi(a)=xa$, where $x$ is given by 18. By Lemmas 3.19 and 3.21 we have $\varphi\in\Gamma(A)$. We are left to prove 22. If $j\neq k$, then
@@
The implications (iv)$\Rightarrow$(iii) and (iii)$\Rightarrow$(i) are obvious. The implication (i)$\Rightarrow$(iv) is a consequence of Lemma 3.22 and the equivalence (ii)$\Leftrightarrow$(iii) is Proposition 3.1. ∎
@@ 4 $\sigma$-matching and interchangeable structures on free non-unital associative algebras
(i)$\Rightarrow$(iii). Let * be $(12)$-matching with $\cdot$. Then for all $a,b,c\in A$ we have
@@
Since $(A,\cdot)$ has no zero divisors, $a\cdot(a*(b\cdot c))=a\cdot((a*b)\cdot c)$ implies $a*(b\cdot c)=(a*b)\cdot c$. Thus, * is totally compatible with $\cdot$ by Remark 1.3(iv).
(ii)$\Rightarrow$(iii). Let * be interchangeable with $\cdot$. Then for all $a,b,c\in A$ we have
@@
As above, $a\cdot((a\cdot b)*c)=a\cdot(a*(b\cdot c))$ yields $(a\cdot b)*c=a*(b\cdot c)$, because $(A,\cdot)$ has no zero divisors. Thus, * is totally compatible with $\cdot$ by Remark 1.3(ii).
The implications (iii)$\Rightarrow$(i) and (iii)$\Rightarrow$(ii) are trivial. ∎
Clearly, $\mathrm{id}$-matching structures on an algebra without zero divisors are not in general totally compatible: as a counterexample it is enough to take a non-commutative associative unital division algebra (see Corollary 2.7).
Free algebras are a classical example of algebras with no zero divisors. In view of Proposition 2.6, we will only be interested in the non-unital case.
@@ 4.1 The free non-unital associative algebra
Given $x,y\in X$, write $x*y=\sum_{u\in X}u\cdot R_{x,y,u}=\sum_{v\in X}L_{x,y,v}\cdot v$ according to 23. Then
@@
for all $x,y,z\in X$. Thus, defining $\star$ to be the restriction of * to $X\times X$, we get 24 and 25 by the associativity of *.
Conversely, let $\star:X\times X\to K\langle X\rangle^{+}$ satisfying 24 and 25. Given $a,b\in X^{+}$, there are unique $x,y\in X$ and $a_{1},b_{1}\in X^{*}$, such that $a=a_{1}x$ and $b=yb_{1}$. Then define
@@
In particular, $x*y=x\star y$ for all $x,y\in X$. The product 26 uniquely extends to a bilinear product * on the whole $K\langle X\rangle^{+}$. Taking additionally $c\in X^{+}$, since $bc=yb_{1}c$, by 26 we have
@@
Thus, we are left to prove that * is associative. Observe from 27, 28 and 24 that 25 is equivalent to
@@
for all $x,y,z\in X$. Let $a=a_{1}x$, $b=yb_{1}$ and $c=zc_{1}$ with $x,y,z\in X$ and $a_{1},b_{1},c_{1}\in X^{*}$. Then by 27 and 28 we have
@@
Case 1. $b_{1}$ is empty. Then using 26, 30, 29 and 28 we have
@@
Case 2. $b_{1}=b_{2}u$ for some $u\in X$ and $b_{2}\in X^{*}$. Then using 26, 30, 27 and 28 we have
@@
We are unable to describe the maps $\star:X\times X\to K\langle X\rangle^{+}$ satisfying 24 and 25 in simpler terms, so let us just give some natural examples. Any associative product $\star:X\times X\to X$, i.e. a semigroup structure on $X$, is of this form. On the other hand, any mutation $x\star y=x\cdot_{p}y$, where $p\in K\langle X\rangle$, is also of this form. We suspect that there also exist examples of more complicated maps $\star$.
@@
It follows from $(x*x)\cdot x=x\cdot(x*x)$ that $p_{x}=q_{x}$, which commutes with $x$ by 34. By Bergman’s centralizer theorem [2] (or directly by induction on the degree of $p_{x}$) one concludes that $p_{x}\in K\langle x\rangle$. Now, 31 gives
@@
so that $p_{x,y}\in K\langle x\rangle\cap K\langle y\rangle=K$. It also follows from 35 that $p_{x,y}=p_{x,v}$ for all $y,v\neq x$. Similarly, 36 gives $p_{x,y}=p_{u,y}$ for all $x,u\neq y$. Let $u\neq v$ and $x\neq y$. If $u\neq y$, then $p_{u,v}=p_{u,y}=p_{x,y}$. If $v\neq x$, then $p_{u,v}=p_{x,v}=p_{x,y}$. If $(u,v)=(y,x)$, then $(x*y)\cdot x=x\cdot(u*v)$ yields $p_{x,y}\cdot y=y\cdot p_{u,v}$, whence $p_{x,y}=p_{u,v}$ because $p_{x,y},p_{u,v}\in K$. Thus, $p_{x,y}$ is a scalar $\lambda\in K$ that does not depend on $x$ and $y$, so the result follows by 33, 35 and 34. ∎
@@
Let $\varphi\in\Gamma(K\langle X\rangle^{+})$ and $*_{\varphi}$ be the corresponding totally compatible product as in Lemma 2.3. Then there exists $\lambda\in K$ such that $\varphi(x\cdot y)=x*_{\varphi}y=\lambda(x\cdot y)$ for all $x,y\in X$ by Lemma 4.4. On the other hand, $\varphi(x\cdot y)=\varphi(x)\cdot y$, whence $\varphi(x)=\lambda x$ for all $x\in X$. Let now $a=xa_{1}\in X^{+}$, where $x\in X$ and $a_{1}\in X^{+}$. Then $\varphi(a)=\varphi(xa_{1})=\varphi(x)\cdot a_{1}=\lambda x\cdot a_{1}=\lambda a$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
@@
In view of Propositions 4.1 and 2.3 we only need to prove (iii)$\Rightarrow$(iv). Let * be totally compatible with $\cdot$. By Lemma 4.4 there exists $\varphi=\lambda\mathrm{id}\in\Gamma(K\langle X\rangle^{+})$, such that $x*y=\varphi(x\cdot y)$ for all $x,y\in X$. Then, given $a=a_{1}x$ and $b=yb_{1}$ with $x,y\in X$ and $a_{1},b_{1}\in X^{*}$, we have
@@ 4.2 The free non-unital commutative associative algebra
It is obvious that a map of the form 37 belongs to $\Gamma(K[X]^{+})$. Conversely, let $\varphi\in\Gamma(K[X]^{+})$.
@@
for all $n\in\mathbb{Z}_{+}$, whence 37 by linearity.
@@
Since $x\neq y$, it follows that $\varphi(x)$ is divisible by $x$ (in $K[X]$), i.e. there exists $p_{x}\in K[X]$ such that $\varphi(x)=xp_{x}$. Then 38 gives $xyp_{y}=xyp_{x}$, so $p_{x}=p_{y}$, which will be denoted by $p$. We thus obtain 37 for $a$ being a variable. Now, for any $x\in X$ and $a_{1}\in K[X]^{+}$ we have $\varphi(xa_{1})=\varphi(x)a_{1}=pxa_{1}$, so 37 holds for monomials of higher degree and thus for all $a\in K[X]^{+}$ by linearity. ∎
@@
(i)$\Rightarrow$(v). Let * be $\mathrm{id}$-matching with $\cdot$. Then by Lemma 4.8 we have
@@
Denoting $p:=x*x$, we get 39 by bilinearity of * and $\cdot$.
(v)$\Rightarrow$(iv). Suppose that * is given by 39. Then $(a*b)*c=a*(b*c)=\frac{a\cdot b\cdot c}{x^{4}}\cdot p^{2}$ and
@@
Observe by Lemma 4.7 that the product 39 from Proposition 4.9(v) is determined by some $\varphi\in\Gamma(K[X]^{+})$ if and only if $\deg(p)\geq 2$.
@@
For all $x\neq y$ in $X$ we have $x\cdot(y*y)=(xy)*y=y\cdot(x*y)$. It follows that $x*y$ is divisible by $x$ and $y*y$ is divisible by $y$ (in $K[X]$). Similarly, changing the brackets in $x*(yx)=x*(xy)$, we get
@@
which shows that $x*y$ is divisible by $y$ (in $K[X]$). Thus, for all $x\in X$ there is $p_{x}\in K[X]$ and for all $x,y\in X$, $x\neq y$, there is $p_{x,y}\in K[X]$ such that
@@
and 40 holds for $x\neq y$ and $\varphi(x)=px$. Moreover, 41 shows that $xyp_{x}=x^{2}yp$, whence
@@
and 40 holds for $x=y$ and the same $\varphi$ too. ∎
@@
(i)$\Rightarrow$(v). Let * be $\mathrm{id}$-matching with $\cdot$. In view of Lemmas 4.8 and 4.11 there exists $\varphi\in\Gamma(K[X]^{+})$ such that
@@
for all $x,y\in X$ and $a_{1},b_{1}\in K[X]$. Then (v) follows by bilinearity.
The implications (v)$\Rightarrow$(iv) and (iv)$\Rightarrow$(i) hold in any associative algebra.
The remaining equivalences are Proposition 4.1. ∎

### v4

@@ 3.2 Algebras with enough idempotents
[…]
@@
The operations ˜13 are clearly well-defined and bilinear. For all $a,b\in A$ and $m\in M$ we have
@@
Since $I_{1}:=\{i\in I\mid\exists k\in I:\ a_{ik}\neq 0\}$ and $I_{2}:=\{j\in I\mid\exists l\in I:\ b_{lj}\neq 0\}$ are finite, then $\operatorname{supp}(amb)$ is finite as a subset of $I_{1}\times I_{2}$ by ˜14. Thus, $AMA\subseteq A$. The converse inclusion is obvious because $a_{i}=e_{i}\cdot e_{i}\cdot a_{i}$ for all $a_{i}\in e_{i}A$, so that $A=A\cdot A\cdot A\subseteq AMA$. ∎
@@
The fact that any $\mathrm{id}$-matching with $\cdot$ product * on $A$ has form ˜16 is a consequence of Lemma˜3.17. Conversely, let * be given by ˜16. Then
@@
Thus, ˜17 holds for all $a\in A$.
Conversely, assume ˜17. Then
@@
We have $e_{k}\cdot(e_{i}*a_{ij})=(e_{k}*e_{i})\cdot a_{ij}=0$ for all $a_{ij}\in e_{i}Ae_{j}$ and $k\neq i$ by Lemma˜3.20, so that
@@
But $e_{i}*a_{ij}=e_{i}*(a_{ij}\cdot e_{j})=(e_{i}\cdot a_{ij})*e_{j}=a_{ij}*e_{j}$, whence ˜19 by ˜20 and 21. ∎
@@
Let $\varphi(a)=xa$, where $x$ is given by ˜18. By Lemmas˜3.19 and 3.21 we have $\varphi\in\Gamma(A)$. We are left to prove ˜22. If $j\neq k$, then
@@
[…]
@@ 4 $\sigma$-matching and interchangeable structures on free non-unital associative algebras
[…]
@@
Since $(A,\cdot)$ has no zero divisors, $a\cdot(a*(b\cdot c))=a\cdot((a*b)\cdot c)$ implies $a*(b\cdot c)=(a*b)\cdot c$. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(iv).
[…]
@@
As above, $a\cdot((a\cdot b)*c)=a\cdot(a*(b\cdot c))$ yields $(a\cdot b)*c=a*(b\cdot c)$, because $(A,\cdot)$ has no zero divisors. Thus, * is totally compatible with $\cdot$ by Remark˜1.3˜(ii).
[…]
Clearly, $\mathrm{id}$-matching structures on an algebra without zero divisors are not in general totally compatible: as a counterexample it is enough to take a non-commutative associative unital division algebra (see Corollary˜2.7).
Free algebras are a classical example of algebras with no zero divisors. In view of Proposition˜2.6, we will only be interested in the non-unital case.
@@ 4.1 The free non-unital associative algebra
Given $x,y\in X$, write $x*y=\sum_{u\in X}u\cdot R_{x,y,u}=\sum_{v\in X}L_{x,y,v}\cdot v$ according to ˜23. Then
@@
for all $x,y,z\in X$. Thus, defining $\star$ to be the restriction of * to $X\times X$, we get ˜24 and 25 by the associativity of *.
Conversely, let $\star:X\times X\to K\langle X\rangle^{+}$ satisfying ˜24 and 25. Given $a,b\in X^{+}$, there are unique $x,y\in X$ and $a_{1},b_{1}\in X^{*}$, such that $a=a_{1}x$ and $b=yb_{1}$. Then define
@@
In particular, $x*y=x\star y$ for all $x,y\in X$. The product ˜26 uniquely extends to a bilinear product * on the whole $K\langle X\rangle^{+}$. Taking additionally $c\in X^{+}$, since $bc=yb_{1}c$, by ˜26 we have
@@
Thus, we are left to prove that * is associative. Observe from ˜27, 28 and 24 that ˜25 is equivalent to
@@
for all $x,y,z\in X$. Let $a=a_{1}x$, $b=yb_{1}$ and $c=zc_{1}$ with $x,y,z\in X$ and $a_{1},b_{1},c_{1}\in X^{*}$. Then by ˜27 and 28 we have
@@
Case 1. $b_{1}$ is empty. Then using ˜26, 30, 29 and 28 we have
@@
Case 2. $b_{1}=b_{2}u$ for some $u\in X$ and $b_{2}\in X^{*}$. Then using ˜26, 30, 27 and 28 we have
@@
We are unable to describe the maps $\star:X\times X\to K\langle X\rangle^{+}$ satisfying ˜24 and 25 in simpler terms, so let us just give some natural examples. Any associative product $\star:X\times X\to X$, i.e. a semigroup structure on $X$, is of this form. On the other hand, any mutation $x\star y=x\cdot_{p}y$, where $p\in K\langle X\rangle$, is also of this form. We suspect that there also exist examples of more complicated maps $\star$.
@@
It follows from $(x*x)\cdot x=x\cdot(x*x)$ that $p_{x}=q_{x}$, which commutes with $x$ by ˜34. By Bergman’s centralizer theorem [2] (or directly by induction on the degree of $p_{x}$) one concludes that $p_{x}\in K\langle x\rangle$. Now, ˜31 gives
@@
so that $p_{x,y}\in K\langle x\rangle\cap K\langle y\rangle=K$. It also follows from ˜35 that $p_{x,y}=p_{x,v}$ for all $y,v\neq x$. Similarly, ˜36 gives $p_{x,y}=p_{u,y}$ for all $x,u\neq y$. Let $u\neq v$ and $x\neq y$. If $u\neq y$, then $p_{u,v}=p_{u,y}=p_{x,y}$. If $v\neq x$, then $p_{u,v}=p_{x,v}=p_{x,y}$. If $(u,v)=(y,x)$, then $(x*y)\cdot x=x\cdot(u*v)$ yields $p_{x,y}\cdot y=y\cdot p_{u,v}$, whence $p_{x,y}=p_{u,v}$ because $p_{x,y},p_{u,v}\in K$. Thus, $p_{x,y}$ is a scalar $\lambda\in K$ that does not depend on $x$ and $y$, so the result follows by ˜33, 35 and 34. ∎
@@
Let $\varphi\in\Gamma(K\langle X\rangle^{+})$ and $_{\varphi}$ be the corresponding totally compatible product as in Lemma˜2.3. Then there exists $\lambda\in K$ such that $\varphi(x\cdot y)=x*_{\varphi}y=\lambda(x\cdot y)$ for all $x,y\in X$ by Lemma˜$4.$ On the other hand, $\varphi(x\cdot y)=\varphi(x)\cdot y$, whence $\varphi(x)=\lambda x$ for all $x\in X$. Let now $a=xa_{1}\in X^{+}$, where $x\in X$ and $a_{1}\in X^{+}$. Then $\varphi(a)=\varphi(xa_{1})=\varphi(x)\cdot a_{1}=\lambda x\cdot a_{1}=\lambda a$. By linearity, $\varphi=\lambda\mathrm{id}$. ∎
@@
[…]
@@ 4.2 The free non-unital commutative associative algebra
It is obvious that a map of the form ˜37 belongs to $\Gamma(K[X]^{+})$. Conversely, let $\varphi\in\Gamma(K[X]^{+})$.
@@
for all $n\in\mathbb{Z}_{+}$, whence ˜37 by linearity.
@@
Since $x\neq y$, it follows that $\varphi(x)$ is divisible by $x$ (in $K[X]$), i.e. there exists $p_{x}\in K[X]$ such that $\varphi(x)=xp_{x}$. Then ˜38 gives $xyp_{y}=xyp_{x}$, so $p_{x}=p_{y}$, which will be denoted by $p$. We thus obtain ˜37 for $a$ being a variable. Now, for any $x\in X$ and $a_{1}\in K[X]^{+}$ we have $\varphi(xa_{1})=\varphi(x)a_{1}=pxa_{1}$, so ˜37 holds for monomials of higher degree and thus for all $a\in K[X]^{+}$ by linearity. ∎
@@
[…]
@@
Denoting p:=xxp:=x*x, we get ˜39 by bilinearity of * and $\cdot$.
[…]
@@
Observe by Lemma˜4.7 that the product ˜39 from Proposition˜4.9˜(v) is determined by some $\varphi\in\Gamma(K[X]^{+})$ if and only if $\deg(p)\geq 2$.
@@
For all $x\neq y$ in $X$ we have $x\cdot(y*y)=(xy)*y=y\cdot(x*y)$. It follows that xyx*y is divisible by $x$ and yyy*y is divisible by $y$ (in $K[X]$). Similarly, changing the brackets in x(yx)=x(xy)x*(yx)=x*(xy), we get
@@
which shows that xyx*y is divisible by $y$ (in $K[X]$). Thus, for all $x\in X$ there is $p_{x}\in K[X]$ and for all $\in X, xyx\neq y$, there is $p_{x,y}\in K[X]$ such that
@@
and ˜40 holds for $x\neq y$ and $\varphi(x)=px$. Moreover, ˜41 shows that $xyp_{x}=x^{2}yp$, whence
@@
and ˜40 holds for $x=y$ and the same $\varphi$ too. ∎
@@
[…]
@@
for all $x,y\in X$ and $a_{1},b_{1}\in K[X]$. Then ˜(v) follows by bilinearity.
[…]
The remaining equivalences are Proposition˜4.1. ∎